# Toward High-Performance LLM Serving: A Simulation-Based Approach for Identifying Optimal Parallelism

### v1

@@ 4.2. Evaluation of APEX’s suggestion
In this experiment, we address the following research question: Can APEX improve LLM serving performance by identifying an optimal parallel execution plan? For evaluation, we design multiple tasks, with each task comprising an LLM, request traces, and an arrival rate. We run APEX simulation for each task and compare the performance of the following three execution plans:
@@
  • Baseline plan: We follow the commonly-used heuristics (Miao et al., 2024; vLLM, v060), which applies tensor parallelism within the same node and pipeline parallelism across nodes. Since we only run on a one-node cluster, we apply tensor parallelism as the baseline plan.
@@
  • APEX Optimal plan: APEX identifies an optimal plan for each task. However, APEX’s search space extends beyond the capabilities of current LLM serving systems, including advanced features like cell-level data parallelism. As a result, an execution plan may not be fully supported by existing LLM serving systems.
@@
  • Feasible Optimal plan: This is the optimal plan identified by APEX under the constraint that only parallelism techniques supported by current LLM serving systems are used. Due to the reduced search space, the feasible optimal plan may achieve lower performance compared to the unconstrained APEX Optimal plan.
While the APEX Optimal plan may not be realizable in current LLM serving systems, we include the results to showcase the potential performance gains that can be achieved by supporting a broader range of parallelism techniques.
@@
We show the experimental results in Table 2. We experiment with the three request traces and four LLMs mentioned in Section 4.1. We also conduct experiments using two different arrival rates for the requests, assuming a Poisson distribution. To explore different quantization formats, we quantize the Mistral-Large model using FP8 for the KV cache, as well as the weights and activations (i.e., W8A8). APEX identifies optimal parallel execution plans for both dense and sparse LLMs (i.e, MoE models), and the Feasible Optimal plans consistently deliver performance improvement over the baseline, achieving up to 2.61$\times$ speedup in terms of end-to-end serving latency. Furthermore, assuming cell-level data parallelism is available, the APEX Optimal Plans can achieve a speedup of up to 4.42$\times$. Cell-level data parallelism is particularly effective when the execution time varies among cells (e.g., MHA and MLP cells). This approach allows a specific cell to be parallelized further using data parallelism without requiring replication of the entire model. In this experiment, we demonstrate APEX is able to identify an optimal parallel execution plan that improves LLM serving performance under various LLMs and request traces.
@@
Figure 5. We compare the speedup predicted by APEX against the actual speedup achieved using vLLM. APEX accurately predicts the speedup of adopting an optimal execution plan over the baseline plan.
@@
APEX identifies an optimal parallel execution plan for each task. The optimal plans consist of various combinations of data, pipeline, and tensor parallelism, effectively balancing the trade-offs of various factors (e.g., compute, memory, network traffic) to outperform the baseline plan, which only relies on tensor parallelism. While the optimal plans are different from task to task, we observe that incorporating data parallelism (DP) often yields performance benefits. Many of the identified optimal plans set the degree of DP to 2 or even 4. Existing LLM serving systems often overlook DP, assuming it is prohibitively memory-intensive. Instead, our results demonstrate that trading memory efficiency for reduced communication overhead can lead to performance improvements. We also observe that the Feasible Optimal plan of Mixtral-8x22B shows marginal improvement over the baseline on the Creation dataset. This is due to the dominating role of memory utilization in this setup, leaving fewer trade-offs available for performance improvement. Specifically, Mixtral-8x22B, with its 141B parameters, heavily utilizes memory resources, while the Creation dataset’s generation-intensive nature further increases memory demands due to extensive KV caching. Under these conditions, the baseline plan, which adopts pure tensor parallelism, performs well owing to its high memory efficiency. Nevertheless, the APEX Optimal plan still manages to deliver a 2.54$\times$ speedup by enabling cell-level data parallelism (DP), which is more memory-efficient than replicating the entire model.
4.3. Evaluation of Simulation Fidelity
In this experiment, we address the following research question: Can APEX provide high-fidelity simulation for LLM serving? While we demonstrate in Table 2 that APEX can identify optimal parallel execution plans that enhance LLM serving performance, it is also essential to verify the accuracy of its simulation results. We evaluate the simulation fidelity from two perspectives. First, under a fixed number of devices, we assess whether APEX can accurately predict the performance improvement of an optimal plan compared to a baseline plan. Second, for a given execution plan, we evaluate whether APEX can reliably predict performance improvements when scaling across different numbers of devices. We show the experimental results in Figure 5 and Figure 6. We set the arrival rate to 0.5 for both experiments as an example; setting the arrival rate of 0.25 yields similar results.
For the first experiment set, we compare the predicted speedup of Feasible Optimal plans with the actual speedup achieved on the device cluster (Figure 5). We do not evaluate Mixtral-8x22B as vLLM currently does not support expert parallelism. APEX achieves accurate predictions, with an average relative error of 9.5%. The largest discrepancies occur with the Qwen-2.5-32B model, the smallest model in our experiment, where the relative error reaches 23%. This is because, for smaller models, the overhead of operations such as RMSNorm becomes relatively significant, and APEX does not account for these operations in its simulations. In contrast, for larger models like Llama-3.1-70B and Mistral-Large (123B), the predictions are more accurate, as the execution time is dominated by attention and feedforward cells, which are well-modeled by APEX. The fidelity for smaller LLMs could be improved by incorporating profiling results for additional operations. However, as APEX primarily focuses on optimizing performance for serving large models, which are more computationally intensive, we leave this improvement for future work. For the second experiment set, we evaluate the time per output token (TPOT) across different numbers of GPUs using tensor parallelism. We show the results in Figure 6. The APEX prediction results are shown as blue solid lines, while the actual results are represented by red dotted lines. The high similarity between the two lines across all cases demonstrates that APEX accurately captures the scalability trend when scaling from 2 GPUs to 8 GPUs across various models and request traces. The estimated TPOT (ms) of APEX is plotted on the left-hand Y-axis, and the actual TPOT (ms) is plotted on the right-hand Y-axis. The actual measured TPOT is consistently higher than the predicted value, as evident from the larger values on the right-hand Y-axis compared to the left-hand Y-axis. This is because APEX primarily focuses on the overhead of key operations, such as attention and feedforward, while omitting the overhead of other operations. Nevertheless, APEX accurately captures the relative performance differences between various parallel execution plans and degrees of parallelism, as demonstrated in Figure 5 and 6. This capability enables APEX to effectively evaluate and compare different execution plans, thereby identifying an optimal parallel execution plan for LLM serving.
@@ 4.4. Evaluation of Efficiency
We evaluate the efficiency of APEX by comparing the actual execution time on the device cluster with the APEX simulation time. Evaluating all parallel execution plans across the setups in Table 2 (i.e., different models, traces, arrival rates) takes approximately 160 hours on 8 H100 GPUs. In contrast, the same evaluation is completed in less than 2.5 hours using APEX on CPU, making it 71$\times$ faster than the actual implementation. From a cost perspective, running the actual implementation would cost approximately \$currency-dollar\\$\$8,889 (based on the Azure NC40ads H100 cluster pricing), whereas the APEX simulation costs only \$currency-dollar\\$\$7.20 (assuming an Azure D64s v6 node). This translates to a 1234.5$\times$ cost reduction, demonstrating the significant efficiency and cost-effectiveness of APEX. Obtaining the operation-level profiling results to set up APEX takes approximately 40 GPU hours. Yet, this is a one-time cost that can be amortized for the same type of hardware device.
@@
We also evaluate the simulation overhead when scaling to larger LLMs on larger device clusters. We use the following example models: Qwen2.5-32B, Llama-3.1-70B, Mistral-Large (123B), and Llama-3.1-405B. We synthesize a trillion-scale model by scaling the Llama-3.1-70B model 16 times; this can be done by modifying values in the LLM configuration file. As shown in Figure 7, APEX demonstrates high scalability, maintaining a similar simulation overhead when scaling from a 32B model to a trillion-scale model. This efficiency is achieved by leveraging the canonical representation in Transformer IR (Section 3.3) and the repetitive structure of the transformer architecture (Section 3.7) to greatly reduce the design space. This scalability highlights APEX’s potential for future applications, as model sizes continue to grow and the cost of actual deployment may become even more expensive.
@@ 4.5. Evaluation of Extensibility
LLMs and their serving systems are evolving rapidly. While APEX supports a wide range of models, devices, etc., the simulator may still become outdated as new advancements emerge. To address this, APEX is designed for extensibility, enabling easy adaptation to support the latest developments. We evaluate APEX’s extensibility by measuring the overheads to implement a new feature. We use two metrics: (1) programming overhead, measured in lines of code required for a feature extension, and (2) implementation time overhead, measured in hours, including time to write and execute the necessary code or scripts. We reported the overhead of various types of extensions in Table 3. Below, we provide details of the extensions implemented for the evaluation.
Extending to New Models APEX can effortlessly support a new LLM without additional programming or implementation time by requiring only a configuration file of the model. However, for LLMs containing unknown transformer cells (Section 3.3), such as a novel feedforward network, additional effort is needed. The primary work involves implementing the Parallel Template (Section 3.4) for the new cells. For evaluation, we measured the effort to support SwiGLU cells (Shazeer, 2020) used in the Llama and Qwen models.
Extending to New Device Cluster APEX can easily adapt to new device clusters by providing the device name, memory capacity, and interconnection network (e.g., NVLink, PCIe). We evaluated the effort required to support a new type of GPU and an in-house AI accelerator. The implementation time overhead mainly involves running profiling scripts, which take 4 to 8 hours depending on the hardware’s performance. While extending to new devices requires a relatively long script execution time, this effort is only necessary when a new device is released.
Extending to New Batching Mechanism APEX can also support new batching mechanisms. While this is not as straightforward as supporting a new model or a new device cluster, it can be done by modifying the Iteration-Level Batching Module 3.6. APEX adopts a vLLM-style (Kwon et al., 2023) batching by default. As an example, we evaluate the overhead of extending to support a Sarathi-Serve-style (Agrawal et al., 2024b) batching, which performs chunk prefilling to better interleave prefill and decode requests. This extension is done by adding a new variable, chunk size, to the batching module and a counter to each request to ensure all prefill chunks are completed before moving to the generation stage.
@@
Extending to New Parallelism The overhead of extending APEX to support new parallelism types depends on the specific parallelism being added. For example, supporting a new type of intra-layer parallelism is analogous to supporting a new transformer cell, as it primarily requires adding a new Parallel Template. Similarly, Fully Sharded Data Parallelism (FSDP) can be integrated into APEX by extending the existing Data Parallelism implementation, incorporating additional collective communication steps, and adjusting memory usage to account for sharded model storage.
In summary, APEX is designed to support a wide range of extensions with minimal programming and implementation overhead, enabling it to stay up-to-date with advancements in LLMs, hardware, and serving systems as they continue to evolve.

### v2

@@ 4.2. Evaluation of APEX’s prediction
In this experiment, we address the following research question: Can APEX improve LLM serving performance by identifying optimal parallel execution plans? For evaluation, we design multiple tasks, with each task comprising an LLM, request trace, and arrival rate. We run APEX simulation for each task and compare the end-to-end serving latency of the following three plans:
@@
1. Baseline plan: We follow the commonly-used heuristics (Miao et al., 2024; vLLM, v060), which applies tensor parallelism within the same node and pipeline parallelism across nodes.
@@
2. APEX Optimal plan: APEX identifies an optimal plan for each task. However, APEX’s search space extends beyond the capabilities of current LLM serving systems, including advanced features like cell-level data parallelism. As a result, an execution plan may not be fully supported by existing LLM serving systems. Still, we include the results to showcase the potential performance gains that can be achieved by supporting a broader range of parallelism techniques.
@@
3. Feasible Optimal plan: This is the optimal plan identified by APEX under the constraint that only parallelism techniques supported by current LLM serving systems are used. Due to the reduced search space, the feasible optimal plan may achieve lower performance compared to the unconstrained APEX Optimal plan.
4.2.1. Single-Node H100 cluster
@@
We first evaluate APEX on a single-node cluster with 8 H100 GPUs and show the experimental results in Table 2. We experiment with the three request traces and three LLMs. We also conduct experiments using two different arrival rates for the requests, assuming a Poisson distribution (Consul and Jain, 1973). To explore different quantization formats, we quantize the Mistral-Large model using FP8 for the KV cache, as well as the weights and activations (i.e., W8A8). APEX identifies optimal parallel execution plans for both dense and sparse LLMs (i.e, MoE models), and the Feasible Optimal plans consistently deliver performance improvement over the baseline, achieving up to 1.75$\times$ speedup in terms of end-to-end serving latency. Furthermore, assuming cell-level data parallelism is supported, APEX Optimal Plans achieve speedups of up to 3.37$\times$. In this experiment, we demonstrate APEX is able to identify an optimal parallel execution plan that improves LLM serving latency under various LLMs and request traces. The identified optimal plans consist of various combinations of data, pipeline, and tensor parallelism, effectively balancing the trade-offs of various factors (e.g., compute, memory, network traffic) to outperform the baseline plan, which only relies on tensor parallelism. While the optimal plans are different from task to task, we observe that incorporating data parallelism (DP) often yields performance benefits. Many of the identified optimal plans set the degree of DP to 2. Existing LLM serving systems often overlook DP, assuming it is prohibitively memory-intensive. Instead, our results demonstrate that trading memory efficiency for reduced communication overhead can lead to performance improvements. We also observe that the Feasible Optimal plan of Mixtral-8x22B shows only marginal improvement over the baseline on the Creation trace. This is due to the dominating role of memory utilization in this setup, leaving fewer trade-offs available for performance improvement. Specifically, Mixtral-8x22B, with its 141B parameters, heavily utilizes memory resources, and the Creation traces’s generation-intensive nature further increases memory demands due to extensive KV caching. Under these conditions, the baseline plan, which adopts pure tensor parallelism, performs well owing to its high memory efficiency. Nevertheless, the APEX Optimal plan still manages to deliver a 1.77$\times$ speedup by enabling cell-level DP.
Table 3. End-to-End latency (sec.). APEX identifies optimal execution plans for different clusters
Model & Cluster Traces Baseline APEX Optimal
Llama-3.1-405B (Multi-Node H100) Summ. 4077.88 2519.12 (1.62$\times$)
Creation 2314.25 1229.14 (1.88$\times$)
Chat 2645.89 1710.10 (1.55$\times$)
Llama-3.1-70B (H200) Summ. 1433.21 897.95 (1.60$\times$)
Creation 1128.66 667.85 (1.69$\times$)
Chat 1189.00 536.74 (2.22$\times$)
Mistral-Large (H200) Summ. 1595.24 1174.79 (1.36$\times$)
Creation 1389.19 776.08 (1.79$\times$)
Chat 1515.41 961.59 (1.56$\times$)
Mixtral-8x22B (H200) Summ. 1540.42 706.06 (2.18$\times$)
Creation 1380.75 993.34 (1.39$\times$)
Chat 1380.76 500.32 (2.76$\times$)
4.2.2. Multi-Node H100 Cluster
Next, we evaluate APEX on a two-node cluster, each node contains 8 NVIDIA H100 SXM GPUs, identical to the setup in Section 4.2.1. We use Llama-3.1-405B, one of the largest publicly available LLMs for evaluation on this cluster. Results are presented in the first row of Table 3. Due to space constraints, we report only the Baseline Plan and APEX Optimal Plan under an arrival rate of 0.5. Experimental results show that APEX consistently identifies execution plans that outperform the baseline, achieving speedups ranging from 1.55$\times$ to 1.88$\times$. These findings demonstrate APEX’s ability to scale effectively to multi-node clusters and extremely large models, continuing to identify performant execution plans under increased system complexity.
4.2.3. H200 Cluster
We also evaluate APEX on a cluster with 8 NVIDIA H200 GPUs. Compared to the H100 (80 GB HBM), the H200 offers the same compute and peak performance but with increased memory capacity (141 GB HBM). This expanded memory enables the H200 cluster to explore a larger design space, allowing for more flexible trade-offs between parallelism strategies. As shown in last three rows of Table 3, APEX finds simulation plans that lead to up to 2.76$\times$ speedup compared to the baseline plan. Notably, APEX recommends different optimal plans for the H200 and H100 clusters under the same experimental setup. For instance, when serving the Summarization trace with the Mistral-Large model, APEX suggests 2-way data parallelism for H100, and 4-way data parallelism plan for H200. While the H200 plan incurs higher memory overhead, it reduces communication overhead, making the trade-off favorable due to the H200’s larger memory capacity.
4.2.4. Optimize Towards Energy Efficiency
For service providers, latency is not the only metric that matters. Rather than minimizing inference time at all costs, providers are typically required to meet specific service-level objectives (SLOs), allowing room to optimize other metrics as long as the SLOs are satisfied. One critical metric is energy consumption, which directly impacts both operational costs and carbon emissions. APEX supports optimizing for energy consumption by profiling not only execution time but also the energy usage of operations. We demonstrate this capability in Table 4, which presents results for serving a Llama-3.1-70B model on a cluster with 8 H100 GPUs. The first rows report latency-optimal plans (i.e., the APEX Optimal Plans in Table 2), which generally yield better energy efficiency than the baseline plans due to shorter runtime. However, latency-optimal plans are not necessarily energy-optimal. We observe that energy efficiency varies across operations of different input sizes, and some operations consume disproportionately more energy; investigating the underlying cause of this phenomenon is beyond the scope of this work, and we instead focus on leveraging the collected profiling results to identify energy-efficient execution plans. APEX evaluates the energy consumption across various parallel execution plans and identifies the energy-optimal plan. Reported in the second rows of Table 4, energy-optimal plans reduce energy consumption by up to 19% compared to latency-optimal plans. For systems with more relaxed SLOs, service providers may further reduce energy consumption by lowering GPU frequency. In the third rows of Table 4, APEX identifies optimal plans under 0.8 GHz of GPU frequency, reducing energy consumption by 45% and 41% on the Summarization and Creation traces, respectively. While this results in increased TTFT (up to 1.9$\times$) and TPOT (up to 1.8$\times$), such increases may still be acceptable in practice, for instance, when the TPOT requirement is 50 ms. These results show that APEX enables service providers to make informed trade-offs between latency and energy efficiency.
Table 4. APEX identifies plans optimized toward energy
Traces
Exe. Plans
(GPU Freq., GHz)
Energy
TTFT
TPOT
Summ. Latency opt. (2.0) 4.095 788.67 32.06
Energy opt. (2.0) 3.293 992.69 26.89
Energy opt. (0.8) 2.265 1485.94 33.62
Creation Latency opt. (2.0) 8.388 44.67 18.83
Energy opt. (2.0) 8.277 54.59 20.63
Energy opt. (0.8) 5.016 75.70 34.59
4.3. Evaluation of Simulation Fidelity
In this experiment, we address the following research question: Can APEX provide high-fidelity simulation for LLM serving? While we demonstrate in Section 4.2 that APEX can identify optimal parallel execution plans that enhance LLM serving performance, it is also essential to verify the fidelity of its simulation results. We evaluate the simulation fidelity from two perspectives. First, under a fixed number of devices, we assess whether APEX can accurately predict the performance improvement of an optimal plan compared to a baseline plan. Second, for a given execution plan, we evaluate whether APEX can reliably predict performance improvements when scaling across different numbers of devices. We show the experimental results in Figure 6 and Figure 7.
@@ 4.3. Evaluation of Simulation Fidelity
Figure 6. Speedup predicted by APEX vs. actual speedup achieved. APEX accurately predicts the speedup of adopting an optimal execution plan over the baseline plan.
For the first experiment set, we compare the predicted speedup of Feasible Optimal plans with the actual speedup achieved on the device cluster (Figure 6). We use vLLM (Kwon et al., 2023) to serve Llama-3.1-70B and Mistral-Large, and use SGLang (Zheng et al., 2024b) for Mixtral-8x22B as vLLM does not support expert parallelism (EP) at the time of our experiments. APEX achieves accurate predictions, with an average relative error of 10.7%. The main discrepancy lies in the predictions for Mixtral-8x22B. APEX estimates that applying EP would outperform tensor parallelism (TP) due to the lower overhead of All-to-All communication in EP, compared to the AllReduce operations required in TP. However, when serving with SGLang, we observe that TP consistently outperforms EP across experiments. A potential reason is that the current implementation of EP may not be as well-optimized as TP, leading to higher-than-expected overheads during execution.
For the second experiment set, we evaluate the time per output token (TPOT) across different numbers of GPUs using tensor parallelism. We show the results in Figure 7. The APEX prediction results are shown as green solid lines, while the actual results are represented by red dotted lines. The high similarity between the two lines across all cases demonstrates that APEX accurately captures the scalability trend when scaling from 2 GPUs to 8 GPUs across various models and request traces. The estimated TPOT (ms) of APEX is plotted on the left-hand Y-axis, and the actual TPOT (ms) is plotted on the right-hand Y-axis. The actual measured TPOT is consistently higher than the predicted value, as evident from the larger values on the right-hand Y-axis compared to the left-hand Y-axis. This is because APEX focuses on the overhead of key operations, such as attention and feedforward, while omitting the overhead of other operations. Nevertheless, APEX accurately captures the relative performance differences between various parallel execution plans and degrees of parallelism, as demonstrated in Figure 6 and 7. This capability enables APEX to effectively evaluate and compare different execution plans, thereby identifying an optimal parallel execution plan for LLM serving.
@@ 4.4. Evaluation of Efficiency
We evaluate the efficiency of APEX by comparing the APEX simulation time with the actual execution time on the device cluster. Evaluating all parallel execution plans across the setups in Table 2 (i.e., different models, traces, arrival rates) takes approximately 160 hours on 8 H100 GPUs. In contrast, the same evaluation is completed in less than 2.5 hours using APEX on a CPU, making it 71$\times$ faster than the actual deployment. From a cost perspective, running the actual implementation would cost approximately \$currency-dollar\\$\$8,889 (based on the Azure NC40ads H100 cluster pricing), whereas the APEX simulation costs only \$currency-dollar\\$\$7.20 (assuming an Azure D64s v6 CPU node). This translates to a 1234.5$\times$ cost reduction, demonstrating the significant time efficiency and cost-effectiveness of APEX. Obtaining the operation-level profiling results to set up APEX takes approximately 40 GPU hours. Yet, this is a one-time cost that can be amortized for the same type of device cluster.
@@
We also evaluate the simulation overhead when scaling to larger LLMs on larger device clusters. We use the following example models: Qwen2.5-32B, Llama-3.1-70B, Mistral-Large, and Llama-3.1-405B. We synthesize a trillion-scale model by scaling the Llama-3.1-70B model 16 times; this is done by modifying values in the LLM configuration file (Huggingface, 2025). As shown in Figure 8, APEX demonstrates high scalability, maintaining a similar simulation overhead when scaling from a 32B to a trillion-scale model. This efficiency is achieved by leveraging the canonical representation of Transformer IR (Section 3.2.1) and the repetitive structure of the transformer architecture to reduce the design space. This scalability highlights APEX’s potential for future applications, as model sizes continue to grow and the cost of actual deployment become more expensive.
@@ 4.5. Evaluation of Extensibility
While APEX supports a wide range of models, devices, etc., the simulator may still become outdated as new advancements emerge. To address this, APEX is designed for extensibility, enabling easy adaptation to support the latest developments. We evaluate APEX’s extensibility by measuring the overheads to implement a new feature. We use two metrics: (1) programming overhead, measured in lines of code required for a feature extension, and (2) implementation time overhead, measured in hours, including time to write and execute the code or scripts. We reported the overhead of various types of extensions in Table 5 and discuss the details below.
Extending to New Models APEX can effortlessly support a new LLM without additional programming or implementation time by requiring only a configuration file (Huggingface, 2025) of the model. However, for LLMs containing unknown transformer cells (Section 3.2.1), such as a novel feedforward network, additional effort is needed. The primary work involves implementing the Parallel Template (Section 3.2.2) for the new cells. For evaluation, we measured the effort to support SwiGLU cells (Shazeer, 2020) used in the Llama-3.1 models.
Extending to New Device Cluster APEX can easily adapt to new device clusters by providing the device name, memory capacity, and interconnection network (e.g., NVLink, PCIe). We evaluated the effort required to support a new GPU model. The implementation time overhead mainly involves running profiling scripts, which take 6 to 8 hours. While extending to new devices requires a relatively long profiling time, this is a one-time effort that can be amortized.
Extending to New Batching Mechanism APEX also supports new batching mechanisms. While this is not as straightforward as supporting a new model or a new device cluster, it can be done by modifying the Batching Module (Section 3.3). APEX adopts contiguous batching (Kwon et al., 2023; Zheng et al., 2024b) by default. We evaluate the overhead of extending to support a Sarathi-Serve-style (Agrawal et al., 2024b) batching, which performs chunked prefill to better interleave prefill and decode requests. This extension is done by adding a new variable, chunk size, to the Batching Module and a counter to each request to ensure all prefill chunks are completed before moving to the decode stage.
@@
Extending to New Parallelism The overhead of extending to new parallelisms depends on the specific parallelism being added. For example, supporting a new intra-layer parallelism is analogous to supporting a new transformer cell, as it requires adding a new Parallel Template. Similarly, Fully Sharded Data Parallelism (FSDP) can be integrated into APEX by extending the existing DP implementation, incorporating additional collective communications and adjusting memory usage for the sharded model storage.